# Generative Active Learning for Image Synthesis Personalization

### v1

@@ Abstract
I Introduction
@@
Recently, generative models, such as large language models (e.g., ChatGPT [1], Llama [2]) and image generation models DALL•E [3], Stable Diffusion [4]), have demonstrated impressive capabilities in producing compelling and diverse results. The key to the success lies in the availability of high-quality training samples on an incredibly large scale. In addition to the real-world datasets that are expensive to collect, numerous studies [5, 6] have demonstrated that incorporating synthetic samples can effectively improve the capability and generalization of models. However, as the number of generated samples can be extensive and of varying quality, a crucial question arises: how can we select the most informative samples with minimal cost for training? This issue has been extensively discussed in the field of active learning, which attempts to maximize a model’s performance while annotating the fewest samples [7]. However, traditional active learning approaches primarily focus on improving discriminative models. The application of active learning in generative models, particularly in utilizing synthetic samples to enhance model performance, remains an open and challenging research area.
@@
In this paper, we present a pilot study on the application of active learning in generative models, specifically focusing on the image synthesis personalization (ISP) [8]. ISP is a representative family of generative tasks that requires the cost-effective selection of synthetic data for training. The learning objective of ISP is to model the user’s “subject of interest” (SoI) based on a limited number of reference images and generate new images that feature the SoI. For instance, in the case of learning from a few images of the user’s pet cat, the trained model should be capable of generating diverse scenes with the cat, such as the cat being situated under the Eiffel Tower or flying alongside Trump, depending on the given prompt [9]. Similarly, when the SoI revolves around a specific style, like Van Gogh paintings or the user’s own artwork, the ISP model should be able to adopt that style and generate new images with the same artistic characteristics [10]. Given the highly specific nature of personal interests, the availability of reference images is often limited. Therefore, selecting good samples from the newly generated images to augment the reference set has proven to be a more practical approach [10]. This can be done in an iterative manner, which aligns well with the framework of active learning.
@@
While the idea of bringing active learning from discriminative models to generative models holds promise, it also presents several challenges. One key challenge is the causal loop in the querying strategy design. In discriminative active learning (DAL), informative samples are selected and queried from a closed set of unlabeled data, typically for tasks like recognizing predefined simple concepts (e.g., dog). This closed set nature makes it feasible to design strategies that compare the information carried by different unlabeled samples (e.g., entropy in uncertainty sampling) so as to prioritize directions in the feature space for querying. In contrast, generative active learning (GAL) faces a scenario where the querying is open to all directions, because the user may combine the SoI with all possible prompts, which can carry much more complex and undetermined semantics. This openness makes the sample-evaluation-based DAL querying strategies infeasible in GAL. It is due to the fact that generated samples are not readily available unless prompts are given, and it is not easy to design prompts before determining the directions to query. This creates a typical causal loop, making it challenging to establish a clear sequence of actions between determining what to generate and knowing which directions to query.
@@
In this paper, we tackle this challenge by transforming the open querying problem into a semi-open one. Our approach involves collecting prompts from existing benchmarks to create a pool of querying intentions. The embeddings of these prompts serve as anchors in the target space, indicating the candidate directions to query and explore. During each iteration of the GAL process, we generate samples using these prompts for evaluation. This semi-open scheme strikes a balance by constraining the candidate directions for querying while allowing enough freedom to explore the target space through the generation of samples. Although this approach provides access to generated samples, the sample-based evaluation commonly used in DAL cannot be directly applied to GAL due to the fundamental differences between discriminative and generative models. Discriminative models learn a single distribution to distinguish simple semantics (e.g., dog), resulting in semantically consistent information carried by positive (negative) samples [11]. On the other hand, generative models focus on generalizing to various mixed semantics (e.g., to generate images not only of the user’s pet dog flying alongside Trump but also of the pet dog chasing butterflies on the beach) [3]. Consequently, generative models need to handle multiple sub-distributions, each modeling a specific combination of semantics. The information carried by samples from different sub-distributions are not consistent, rendering sample-based evaluation infeasible. To address this issue, we propose a distribution-based querying strategy that adapts the classical Uncertainty Sampling [7, 12] to the new generative scenario. This strategy takes into account the distributional aspects of generative models and provides a more suitable framework for querying and evaluating samples in GAL.
@@
Another challenge is the exploitation-exploration dilemma [13, 14]. In DAL, the collected samples from different iterations are accumulated for training, and the learned distribution or decision boundary may gradually shift from the samples collected in the early iterations. This is generally not a problem as long as it benefits the classifier’s performance. In contrast, in GAL, the fidelity to the reference images is of great importance which pushes the generated samples towards the references. Additionally, samples generated in the early iterations have been shown to have a higher likelihood of fulfilling the fidelity criteria compared to later iterations, and thus should be exploited as new references with greater attention. However, the generated samples cannot be too close to the references, otherwise, this causes over-fitting. Meanwhile, the generated samples need to be generalized to a certain target direction indicated by corresponding prompt, which attracts them to move toward the target direction against the references. The GAL process needs to learn how to navigate this balance between adhering to the references and exploring new directions. We propose a balancing scheme that evaluates the importance of references, thereby allowing us to weigh the contributions of different iterations.
@@
II Related Work
II-A Active Learning
@@
Active learning is a subfield of machine learning, which aims to find an optimal querying strategy to maximize model performance with the fewest labeling cost. The most common strategies include uncertainty sampling [7, 12], query by committee [15], and representation-based sampling [16, 17], etc. The rationale behind is to provide the most valuable samples to learn a better decision boundary. However, acquiring real-world datasets still poses challenges in certain scenarios, such as few-shot learning. To address this issue, some researchers have explored the use of generative networks for data augmentation. For example, GAAL [18] first introduced GAN [19] to generate training samples. But this random generation does not guarantee more informative samples compared to the original dataset. In contrast, BGADL [20] jointly trained a generative network and a classifier so as to generate samples in disagreement regions [7]. Subsequent approaches, such as VAAL [21] and TA-VAAL [22] employed adversarial training for data augmentation to improve the feature representation. It is important to note that while these works have explored the use of generative models, their primary focus is on improving the discriminative model’s ability.
II-B Personalized Content Generation
@@
Text-to-image synthesis has earned significant attention for its potential applications in content creation, virtual reality, and computer graphics. Impressive works such as DALL•E [3], Stable Diffusion [4], Imagen [23], have shown immense potential to generate compelling and diverse images. As an application of image generation, personalized image synthesis offers user an opportunity to create customized object or style that is difficult to generate using pre-trained models. To accomplish content personalization, some studies [24, 25, 26] have concentrated on training a unified model capable of personalizing any input image. However, these approaches often struggle to achieve satisfactory fidelity with the references. In contrast, other research studies [9, 27] adopt a fine-tuning approach on pre-trained models for each reference group. These methods first predefine a pseudo word to represent the target visual concepts and then fine-tune models using provided references, aiming to establish a mapping between the predefined words and the corresponding visual elements. In particular, DreamBooth [9] retrains the entire diffusion model and incorporates a prepared regularization dataset to alleviate the overfitting problem. Following this regularization approach, other works focus on enhancing different aspects of personalized image synthesis, like training acceleration [27] and multiple concepts composition [28, 29, 30]. SVDiff [31], on the other hand, applies image stitching techniques to expand the training samples, but does not explore the utilization of generated samples. Different from DreamBooth, Textual Inversion [32] aims to find an optimal token embedding to reconstruct the training images without additional regularization samples. Although additional training samples are adopted in the training process, no generated samples are involved in these studies.
II-C Personalized Style Generation
Style generation is one of the notable advancements in the field of image synthesis. Style transfer [33, 34, 35] aims to transform the visual style of a given image to another input image while preserving its contents. However, these methods do not offer the chance to generate images based on text prompts. Meanwhile, another line of research focuses on personalized style generation, which aims to reverse visual styles on textual descriptions. A recent study, StyleDrop [10], introduces a parameter-efficient fine-tuning method and an iterative training framework with feedback to facilitate style recreation. Specifically, preset prompts are used to generate images and these images are then subject to user filtering, where users will identify high-quality images that can be used for further training. While this approach leverages human feedback to enhance model performance, the need for human inspection and the equal weighting of selected samples pose limitations. In this paper, we propose methods that effectively alleviate the burden on human resources through active learning and reduce selection bias by balancing the importance of synthetic and real samples.
@@ II-C Personalized Style Generation
  Input: anchor embedding sets $\mathcal{A}$, reference images $\mathbf{x}$, subject of interest $\mathbf{e}^{*}$, non-SoI $\bar{\mathbf{e}}^{*}$, pre-trained model $f_{\mathbf{\theta}}$
@@
           Verify whether $\mathbf{I}_{ij}$ is overfitted by 10
@@
        Calculate $\Omega(\mathbf{a}_{i})$ according to 6
@@
     Update openness score according to 11
  until Stopping criterion is met according to 12
III Method
In this section, we introduce our implementation of generative active learning for image synthesis personalization. The algorithm, along with its corresponding pseudo-code, is depicted in 1.
III-A Preliminaries for Image Synthesis Personalization
The current state-of-the-art methods for Image Synthesis Personalization (ISP) are all based on diffusion models [36, 37]. What sets diffusion models apart is their “generate-by-denoise” approach. During training, a text-image pair is used, and the process begins by iteratively adding noise to the image $\mathbf{x}$ according to the Markov chain, resulting in a noisy image ${\mathbf{x}_{t}}$. The noisy image is then combined with the text embedding $\mathbf{e}$ to create a new noisy image embedded with the text semantics, denoted as $\mathbf{x}_{t}\circ\mathbf{e}$. Learning then proceeds to denoise this image and reconstruct the original image $\mathbf{x}$, which is represented as
$\hat{\mathbf{x}}=f_{\mathbf{\theta}}(\mathbf{x}_{t}\circ\mathbf{e})$ (1)
@@ III-A Preliminaries for Image Synthesis Personalization
$L_{rec}=\mathbb{E}\left[w_{t}\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}\right]$ (2)
where $w_{t}$ represents a time-step dependent weight. In the generation phase, the embedding of a user prompt $\tilde{\mathbf{e}}$ is then fused with a random noise $\mathbf{\epsilon}$ to generate the image $\tilde{\mathbf{x}}=f_{\mathbf{\theta}}(\mathbf{\epsilon}\circ\tilde{\mathbf{e}})$ that align with the semantics of interest.
@@
Figure 1: Examples of overfitted and well-aligned generations. The model has to exclude the non-SoI for successful generations.
III-B Direction-based Uncertainty Sampling
@@
$\mathbf{e}^{*}\oplus\mathcal{A}=\{\mathbf{e}^{*}\oplus\mathbf{a}_{i}\},i\in\mathbb{N}.$ (3)
where $\oplus$ is a model-dependent operator. In the majority of models that employ CLIP [38] as the text encoder, the operation $\oplus$ is typically implemented by directly inputting the embeddings as a sequence.
In each iteration of the generative active learning (GAL), we generate $m$ samples for each anchor prompt. We need to initiate the next round of GAL by selecting informative ones from the generated samples as new references. However, as discussed, conventional sample-based querying is infeasible in GAL, because evaluating performance on individual samples lacks of global perspective to measure the model’s generalizability. Additionally, relying solely on generalizability to build a metric is challenging because higher generalizability may indicate well-explored directions, where samples would not provide novel information for improving the model. This is similar to the situation in Discriminative Active Learning (DAL), where including samples from well-classified locations does not contribute to performance improvement and instead hinders exploration. A popular solution is Uncertainty Sampling [7], which selects samples from areas where the model exhibits uncertainty. In the context of GAL, we can adapt this idea to identify directions where the quality of model-generated samples lies between well-generalized and overfitted. Let $\mathbf{I}_{ij},j\in[1,m]$ denote an image generated for the $i^{th}$ anchor direction $\mathbf{a}_{i}$ as
$\mathbf{I}_{ij}=f_{\mathbf{\theta}}\left(\mathbf{\epsilon}\circ(\mathbf{e}^{*}\oplus\mathbf{a}_{i})\right)$ (4)
@@
$\Omega(\mathbf{a}_{i})$ $=-\left[(1-\beta_{i})\log(1-\beta_{i})+\beta_{i}\log\beta_{i}\right]$ (6)
$\beta_{i}$ $=\frac{\sum_{j=1}^{m}\Phi(\mathbf{I}_{ij})}{m}.$ (7)
@@
In DAL, the learning employs human annotators as oracles. However, due to the computational expense of current diffusion models, it becomes impractical for human annotators to wait for the results of each iteration, resulting in significant delays. Hiring human annotators as oracles can be extremely costly, which might be one of the reasons why successful ISP models using generated results as argumentation are predominantly developed by large companies like Google [10, 39], who can afford such expenses. In our study, we found the oracle function $\Phi(\mathbf{I}_{ij})$ can be estimated by evaluating the generated image $\mathbf{I}_{ij}$’s fidelity to both the anchor direction and irrelevant semantics in the reference prompt. This observation stems from the fact that the reference prompt consists of two components: the SoI and non-SoI semantics. Most previous studies focus on the fidelity to the SoI semantics, while the non-SoI semantics are not fully leveraged. These non-SoI semantics can be considered distractor semantics that the generated images should avoid, similar to negative labels in discriminative models. One such example can be found from 1, in which the SoI is the drawing style while the non-SoI is the concept cat. The overfitted samples are those failed to disentangle the cat from the generation. Therefore, we propose a straightforward metric to simulate the oracle function. Let $\bar{\mathbf{e}}^{*}$ denote the non-SoI embedding, the function is written as
$\Phi(\mathbf{I}_{ij})=$$\left\{\begin{array}[]{cc}1,&sim(\mathbf{I}_{ij},\mathbf{a}_{i})\leq sim(\mathbf{I}_{ij},\bar{\mathbf{e}}^{*})\\ 0,&sim(\mathbf{I}_{ij},\mathbf{a}_{i})>sim(\mathbf{I}_{ij},\bar{\mathbf{e}}^{*})\\ \end{array}\right.$ (10)
@@
where the $sim()$ is a fidelity metric of an image to a semantics. In this study, we simply adopt the CLIP similarity [38].
@@
With all necessary components built, the querying can then be conducted by evaluating all the anchor directions and select the ones with top-$k$ uncertainty scores (using 6). For each direction, we choose $n$ generated images with the highest $sim(\mathbf{I}_{ij},\mathbf{a}_{i})$ scores (indication of the faithfulness to the directions) as new reference images.
III-C Balancing the Exploitation and Exploration

### v2

@@ Abstract.
Figure 1. Given a few images of the subject of interest, the proposed method is capable of generating diverse personalized images in different contexts, such as moon, throne, rococo, etc.
1. Introduction
@@
Recently, generative models, such as large language models (e.g., ChatGPT (Ouyang et al., 2022), Llama (Touvron et al., 2023)) and image generation models DALL•E (Ramesh et al., 2021), Stable Diffusion (Rombach et al., 2022)), have demonstrated impressive capabilities in producing compelling and diverse results. The key to the success lies in the availability of high-quality training samples on an incredibly large scale. In addition to the real-world datasets that are expensive to collect, numerous studies (Nikolenko, 2021; Azizi et al., 2023) have demonstrated that incorporating synthetic samples can effectively improve the capability and generalization of models. However, as the number of generated samples can be extensive and of varying quality, a crucial question arises: how can we select the most informative samples with minimal cost for training? This issue has been extensively discussed in the field of active learning, which attempts to maximize a model’s performance while annotating the fewest samples (Tharwat and Schenck, 2023). However, traditional active learning approaches primarily focus on improving discriminative models. The application of active learning in generative models, particularly in utilizing synthetic samples to enhance model performance, remains an open and challenging research area.
@@
In this paper, we present a pilot study on the application of active learning in generative models, specifically focusing on the image synthesis personalization (ISP) (Po et al., 2023). ISP is a representative family of generative tasks that requires the cost-effective selection of synthetic data for training. The learning objective of ISP is to model the user’s “subject of interest” (SoI) based on a limited number of reference images and generate new images that feature the SoI. For instance, in the case of learning from a few images of the user’s pet cat, as illustrated in Figure 1, the trained model should be capable of generating diverse scenes with the cat, such as the cat on the moon or sitting on the Iron Throne, depending on the given prompt (Ruiz et al., 2023). Similarly, when the SoI revolves around a specific style, like Van Gogh paintings or the user’s own artwork, the ISP model should be able to adopt that style and generate new images with the same artistic characteristics (Sohn et al., 2023). Given the highly specific nature of personal interests, the availability of reference images is often limited. Therefore, selecting good samples from the newly generated images to augment the reference set has proven to be a more practical approach (Sohn et al., 2023). This can be done in an iterative manner, which aligns well with the framework of active learning.
@@
While the idea of bringing active learning from discriminative models to generative models holds promise, it also presents several challenges. One key challenge is the causal loop in the querying strategy design. In discriminative active learning (DAL), informative samples are selected and queried from a closed set of unlabeled data, typically for tasks like recognizing predefined simple concepts (e.g., dog). This closed-set nature makes it feasible to design strategies that compare the information carried by different unlabeled samples (e.g., entropy in uncertainty sampling) so as to prioritize directions in the feature space for querying. In contrast, generative active learning (GAL) faces a scenario where the querying is open to all directions, because the user may combine the SoI with all possible prompts, which can carry much more complex and undetermined semantics. This openness makes the sample-evaluation-based DAL querying strategies infeasible in GAL. This is because generated samples are not readily available unless prompts are given, and it is not easy to design prompts before determining the directions to query. This creates a typical causal loop, making it challenging to establish a clear sequence of actions between determining what to generate and knowing which directions to query.
@@
In this paper, we tackle this challenge by transforming the open querying problem into a semi-open one. Our approach involves collecting prompts to create a pool of querying intentions. The prompt embeddings serve as anchors in the target space, indicating the candidate directions to query and explore. During each iteration of the GAL process, we generate samples using these prompts for evaluation. This semi-open scheme strikes a balance by constraining the candidate directions for querying while allowing enough freedom to explore the target space through the generation of samples. Although this approach provides access to generated samples, the sample-based evaluation commonly used in DAL cannot be directly applied to GAL due to the fundamental differences between discriminative and generative models. Discriminative models learn a single distribution to distinguish simple semantics (e.g., dog), resulting in semantically consistent information carried by positive (negative) samples (Rish et al., 2001). However, generative models focus on generalizing to various mixed semantics (e.g., to generate images not only of the user’s pet dog in a forest but also of the dog on Tokyo street) (Ramesh et al., 2021). Consequently, generative models need to handle multiple sub-distributions, each modeling a specific combination of semantics. The information carried by samples from different sub-distributions are not consistent, rendering sample-based evaluation infeasible. To address this issue, we propose a distribution-based querying strategy that adapts the classical Uncertainty Sampling (Tharwat and Schenck, 2023; Wang and Shang, 2014) to the new generative scenario. It considers the distributional aspects of generative models and provides a more suitable framework for querying and evaluating samples in GAL.
@@
Another challenge is the exploitation-exploration dilemma (Wei and Yang, 2011). In DAL, the collected samples from different iterations are accumulated for training, and the learned distribution or decision boundary may gradually shift from the samples collected in the early iterations. This is generally not a problem as long as it benefits the classifier’s performance. In contrast, in GAL, the fidelity to the reference images is of great importance which pushes the generated samples towards the references. Additionally, samples generated in the early iterations have been shown to have a higher likelihood of fulfilling the fidelity criteria compared to later iterations, and thus should be exploited as new references with greater attention. However, the generated samples cannot be too close to the references, otherwise, this causes over-fitting. Meanwhile, the generated samples need to be generalized to a certain target direction indicated by corresponding prompt, which attracts them to move toward the target direction against the references. The GAL process needs to learn how to navigate this balance between adhering to the references and exploring new directions. We propose a balancing scheme that evaluates the importance of references, thereby allowing us to weigh the contributions of different iterations.
@@
2. Related Work
2.1. Active Learning
@@
Active learning is a subfield of machine learning, which aims to find an optimal querying strategy to maximize model performance with the fewest labeling cost. The most common strategies include uncertainty sampling (Tharwat and Schenck, 2023; Wang and Shang, 2014), query by committee (Seung et al., 1992), and representation-based sampling (Sener and Savarese, 2017; Geifman and El-Yaniv, 2017), etc. The rationale behind is to provide the most valuable samples to learn a better decision boundary. However, acquiring real-world datasets still poses challenges in certain scenarios, such as few-shot learning. To address this issue, the use of generative networks for data augmentation has been investigated. For example, GAAL (Zhu and Bento, 2017) first introduced GAN (Goodfellow et al., 2014) to generate training samples. However, this random generation does not guarantee more informative samples compared to the original dataset. In contrast, BGADL (Tran et al., 2019) jointly trained a generative network and a classifier so as to generate samples in disagreement regions (Tharwat and Schenck, 2023). Subsequent approaches, such as VAAL (Sinha et al., 2019) and TA-VAAL (Kim et al., 2021) employed adversarial training for data augmentation to improve the feature representation. It is important to note that while these works have explored the use of generative models, their primary focus is on improving the discriminative model’s ability.
2.2. Personalized Content Generation
@@
Text-to-image synthesis has earned significant attention for its potential applications in content creation, virtual reality, and computer graphics. Impressive works such as DALL•E (Ramesh et al., 2021), Stable Diffusion (Rombach et al., 2022), Imagen (Saharia et al., 2022), have shown immense potential to generate compelling and diverse images. As an application of image generation, personalized image synthesis offers user an opportunity to create customized object or style that is difficult to generate using pre-trained models. To accomplish content personalization, some studies (Wei et al., 2023; Ye et al., 2023; Li et al., 2023; Chen et al., 2023a) have concentrated on training a unified model capable of personalizing any input image. However, these approaches struggle to perform satisfactory fidelity with the references. In contrast, other research studies (Ruiz et al., 2023; Kumari et al., 2023; Chen et al., 2023b; Arar et al., 2023) enhance subject appearance preservation by adopting fine-tuning approach on pre-trained models for each reference group. In particular, Textual Inversion (Gal et al., 2022) aims to find an optimal token embedding to reconstruct the training images without additional regularization samples. DreamBooth (Ruiz et al., 2023) retrains the entire diffusion model and incorporates a prepared regularization dataset to alleviate the overfitting problem. Following this training framework and regularization approach, other works focus on enhancing different aspects of personalized image synthesis, like training acceleration (Kumari et al., 2023) and multiple concepts composition (Tewel et al., 2023; Liu et al., 2023; Zhang et al., 2024). As for expanding training samples, SVDiff (Han et al., 2023) applies image stitching techniques, but does not explore the use of generated samples. In summary, although additional training samples are adopted in the training process, no generated samples are involved in these studies.
@@
  Input: anchor embedding set $\mathcal{A}$, reference images $\mathbf{x}$, subject of interest $\mathbf{e}^{*}$, non-SoI $\bar{\mathbf{e}}^{*}$, pre-trained model $f_{\mathbf{\theta}}$, number of synthetic samples per prompt $m$
@@
           Verify whether $\mathbf{I}_{ij}$ is overfitted by Equation 10
@@
        Calculate $\Omega(\mathbf{a}_{i})$ according to Equation 6
@@
     Update openness score according to Equation 11
  until Stopping criterion is met according to Equation 12
2.3. Personalized Style Generation
Style generation is one of the notable advancements in the field of image synthesis. Style transfer (Gatys et al., 2016; Zhang et al., 2023; Wang et al., 2023) aims to transform the visual style of a given image to another input image while preserving its contents. However, these methods do not offer the chance to generate images based on text prompts. Meanwhile, another line of research focuses on personalized style generation, which aims to reverse visual styles on textual descriptions. A recent study, StyleDrop (Sohn et al., 2023), introduces a parameter-efficient fine-tuning method and an iterative training framework with feedback to facilitate style recreation. Specifically, preset prompts are used to generate images and these images are then subject to user filtering, where users will identify high-quality images that can be used for further training. While this approach leverages human feedback to enhance model performance, the need for human inspection and the equal weighting of selected samples pose limitations. In this paper, we propose methods that effectively alleviate the burden on human resources through active learning and reduce selection bias by balancing the importance of synthetic and real samples.
3. Method
In this section, we introduce our implementation of generative active learning for image synthesis personalization. The algorithm, along with its pseudo-code, is depicted in Algorithm 1.
3.1. Preliminaries for Image Synthesis Personalization
The current state-of-the-art methods for Image Synthesis Personalization (ISP) are all based on diffusion models (Ho et al., 2020; Song et al., 2020). What sets diffusion models apart is their “generate-by-denoise” approach. During training, a text-image pair is used, and the process begins by iteratively adding noise to the image $\mathbf{x}$ according to the Markov chain, resulting in a noisy image ${\mathbf{x}_{t}}$. The noisy image is then combined with the text embedding $\mathbf{e}$ to create a new noisy image embedded with the text semantics, denoted as $\mathbf{x}_{t}\circ\mathbf{e}$. Learning then proceeds to denoise this image and reconstruct the original image $\mathbf{x}$, which is represented as
(1) $\hat{\mathbf{x}}=f_{\mathbf{\theta}}(\mathbf{x}_{t}\circ\mathbf{e})$
@@ 3.1. Preliminaries for Image Synthesis Personalization
(2) $L_{rec}=\mathbb{E}\left[w_{t}\|\hat{\mathbf{x}}-\mathbf{x}\|_{2}^{2}\right]$
where $w_{t}$ is a time-dependent weight. During the inference, the prompt embedding $\tilde{\mathbf{e}}$ is then fused with a random noise $\mathbf{\epsilon}$ to generate the image $\tilde{\mathbf{x}}=f_{\mathbf{\theta}}(\mathbf{\epsilon}\circ\tilde{\mathbf{e}})$ that aligns with the semantics of interest.
@@
Figure 2. Overfitted and well-aligned generations. The model has to exclude the non-SoI for successful generations.
3.2. Direction-based Uncertainty Sampling
@@
(3) $\mathbf{e}^{*}\oplus\mathcal{A}=\{\mathbf{e}^{*}\oplus\mathbf{a}_{i}\},i\in\mathbb{N}.$
where $\oplus$ is a model-dependent operator, which is typically implemented by directly inputting the embeddings as a sequence.
In each iteration of the generative active learning (GAL), we generate $m$ samples for each anchor prompt. We need to initiate the next round of GAL by selecting informative ones from the generated samples as new references. However, as discussed, conventional sample-based querying is infeasible in GAL, because evaluating performance on individual samples lacks of global perspective to measure the model’s generalizability. Additionally, relying solely on generalizability to build a metric is challenging because higher generalizability may indicate well-explored directions, where samples would not provide novel information for improving the model. This is similar to the situation in Discriminative Active Learning (DAL), where including samples from well-classified locations does not contribute to performance improvement and instead hinders exploration. A popular solution is Uncertainty Sampling (Tharwat and Schenck, 2023), which selects samples from areas where the model exhibits uncertainty. In the context of GAL, we can adapt this idea to identify directions where the quality of model-generated samples lies between well-generalized and overfitted. Let $\mathbf{I}_{ij},j\in[1,m]$ denote an image generated for the $i^{th}$ anchor direction $\mathbf{a}_{i}$ as
(4) $\mathbf{I}_{ij}=f_{\mathbf{\theta}}\left(\mathbf{\epsilon}\circ(\mathbf{e}^{*}\oplus\mathbf{a}_{i})\right)$
@@
(6) $\Omega(\mathbf{a}_{i})$ $=-\left[(1-\beta_{i})\log(1-\beta_{i})+\beta_{i}\log\beta_{i}\right]$
(7) $\beta_{i}$ $=\frac{\sum_{j=1}^{m}\Phi(\mathbf{I}_{ij})}{m}.$
@@
In DAL, the learning employs human annotators as oracles. However, due to the computational expense of current diffusion models, it becomes impractical for human annotators to wait for the results of each iteration, resulting in significant delays. Hiring human annotators as oracles can be extremely costly, which might be one of the reasons why successful ISP models using generated results as argumentation are predominantly developed by large companies like Google (Sohn et al., 2023; Chen et al., 2023a), who can afford such expenses. In our study, we found the oracle function $\Phi(\mathbf{I}_{ij})$ can be estimated by evaluating the generated image $\mathbf{I}_{ij}$’s fidelity to both the anchor direction and irrelevant semantics in the reference prompt. This observation stems from the fact that the reference prompt consists of two components: the SoI and non-SoI semantics. Most previous studies focus on the fidelity to the SoI semantics, while the non-SoI semantics are not fully leveraged. These non-SoI semantics can be considered distractor semantics that the generated images should avoid, similar to negative labels in discriminative models. One such example can be found in Figure 2, in which the SoI is the drawing style while the non-SoI is the concept cat. The overfitted samples are those failed to disentangle the cat from the generation. Therefore, we propose a straightforward metric to simulate the oracle function. Let $\bar{\mathbf{e}}^{*}$ denote the non-SoI embedding, the function is written as
(10) $\Phi(\mathbf{I}_{ij})=$$\left\{\begin{array}[]{cc}1,&sim(\mathbf{I}_{ij},\mathbf{a}_{i})\leq sim(\mathbf{I}_{ij},\bar{\mathbf{e}}^{*})\\ 0,&sim(\mathbf{I}_{ij},\mathbf{a}_{i})>sim(\mathbf{I}_{ij},\bar{\mathbf{e}}^{*})\\ \end{array}\right.$
@@
where the $sim()$ is a fidelity metric of an image to a semantics. In this study, we simply adopt the CLIP similarity (Radford et al., 2021).
@@
With all necessary components built, the querying can then be conducted by evaluating all the anchor directions and selecting the ones with top-$k$ uncertainty scores (using Equation 6). For each direction, we choose the generated image with the highest $sim(\mathbf{I}_{ij},\mathbf{a}_{i})$ score (indication of the faithfulness to the direction) as a new reference image.
3.3. Balancing the Exploitation and Exploration